# Perception, Reason, Think, and Plan: A Survey on Large Multimodal Reasoning Models

### v1

@@ 4.3 Technical Prospects
Takeaways: Native Large Multimodal Reasoning Model (LMRMs) In this section, we examined the latest large multimodal model (e.g., O3 and O4-mini) and their performance on challenging tasks and benchmarks. We then presented the future trajectory for native multimodal large models in terms of capability scope and level, including omnimodal perception and understanding, multimodal interactive generative reasoning, and intelligent agent behavior. To realize this vision, we discussed approaches related to unified perception, learning methods, and data synthesis. We hope that native LMRMs will achieve comprehensive perception, precise understanding, and deep reasoning as a paradigm shift in machine intelligence.
@@ 5 Dataset and Benchmark
Table 9: Overview of Multimodal Benchmarks and Datasets (Training), categorized by task: Understanding (Visual-centric, Audio-centric), Generation (Cross-modal, Joint Multimodal), Reasoning (General Visual, Domain-Specific), and Planning (GUI, Embodied & Simulated Environments). These benchmarks often require short or long reasoning for successful task completion, e.g., challenging visual and audio generation.
@@
Seed-Bench (Li et al., 2023c), MME-RealWorld (Zhang et al., 2024f) MS-COCO (Lin et al., 2014b), Virpt (Yang et al., 2024a)
@@
MMStar (Chen et al., 2024f), TRIG-Bench (Li et al., 2025d) VideoCC3M (Nagrani et al., 2022), FILIP (Yao et al., 2021)
@@
TikTalkCoref (Li et al., 2025f), AgMMU (Gauba et al., 2025) TikTalkCoref (Li et al., 2025f), MRES-32M (Liu et al., 2025a)
@@
RefCOCOm (Liu et al., 2025a)
SBVQA (Alasmary & Al-Ahmadi, 2023), H2VU-Benchmark (Wu et al., 2025)
4D-Bench (Zhu et al., 2025), V2P-Bench (Zhao et al., 2025e)
RSMMVP (Adejumo et al., 2025), HIS-Bench (Zhao et al., 2025b)
MMLA (Zhang et al., 2025b), SARLANG-1M (Wei et al., 2025b)
@@
Air-Bench (Yang et al., 2024d), MMAU (Sakshi et al., 2024) CoVoST2 (Wang et al., 2020), SIFT-50M (Pandey et al., 2025)
@@
MusicNet (Thickstun et al., 2017), AVE-PM (Liu et al., 2025b) ClothoAQA (Lipping et al., 2022), MusicNet (Thickstun et al., 2017)
ACVUBench (Yang et al., 2025f) NSynth (Engel et al., 2017), MusicCaps (Agostinelli et al., 2023)
Multimodal Generation Cross-modal Generation GenEval (Ghosh et al., 2023), T2I-CompBench++ (Huang et al., 2025a) MS-COCO (Lin et al., 2014b), Flickr30k (Plummer et al., 2017)
@@
MJHQ-30K (Li et al., 2024d), VBench 2.0  (Zheng et al., 2025b) UniSim (Yang et al., 2024e), VidProM (Wang & Yang, 2024)
@@
LMM4LMM (Wang et al., 2025e), RISEBench (Zhao et al., 2025d) HQ-Edit (Hui et al., 2024), InstructPix2Pix (Brooks et al., 2023)
WikiVideo (Martin et al., 2025), OmniDiff (Liu et al., 2025c) MagicBrush (Zhang et al., 2023d)
@@
Joint Multimodal Generation MM-Interleaved (Tian et al., 2024), ANOLE (Chern et al., 2024) DreamLLM (Dong et al., 2023), SEED-Story (Yang et al., 2024f)
InterleavedEval (Liu et al., 2024c), OpenLEAF (An et al., 2024) NextGPT (Wu et al., 2024a), DreamFactory (Xie et al., 2024b)
OpenING (Zhou et al., 2024a), M2RAG (Ma et al., 2025c) DreamRunner (Wang et al., 2024o), EVA (Chi et al., 2024)
SEED-Bench (Li et al., 2023c), SEED-Bench-2 (Li et al., 2023b)
MME-Unify (Xie et al., 2025b)
@@
PhysBench (Chow et al., 2025), MMBench (Liu et al., 2024f) MMPR (Wang et al., 2024g), ChartQA (Masry et al., 2022)
@@
VCRBench (Qi et al., 2025), VisualPuzzles (Song et al., 2025d) ShareGPT4oReasoning (Zhang et al., 2024e), R1-Onevision (Yang et al., 2025e)
IV-Bench (Ma et al., 2025a), VisuLogic (Xu et al., 2025d) Video-R1-data (Feng et al., 2025b), Visual-CoT (Shao et al., 2024)
FG-BMK (Yu et al., 2025), Video-MMLU (Song et al., 2025a)
@@
HAVEN (Gao et al., 2025a), MAGIC-VQA (Yang et al., 2025c)
@@
FortisAVQA (Ma et al., 2025b), VideoVista-CulturalLingo (Chen et al., 2025e)
@@
ScienceQA (Lu et al., 2022), ChartQA ((Masry et al., 2022))
@@
VCBench (Wang et al., 2025h), VCT (Götting et al., 2025)
3MDBench (Sviridov et al., 2025), PuzzleBench (Zhang et al., 2025e)
ColorBench (Liang et al., 2025b), VisualPuzzles (Song et al., 2025d)
@@
EIBench (Lin et al., 2025b), XLRS-Bench (Wang et al., 2025b)
STI-Bench (Li et al., 2025i), EgoToM (Li et al., 2025k)
DomainCQA (Zhong et al., 2025),
MMCR-Bench (Yan et al., 2025a), Misleading ChartQA (Chen et al., 2025g)
@@
MicroVQA (Burgess et al., 2025), MPBench (Xu et al., 2025e)
Open3DVQA (Zhan et al., 2025), ProBench (Yang et al., 2025d)
Chart-HQA (Chen et al., 2025c), MMSciBench (Ye et al., 2025)
@@
SWE-BENCH M (Yang et al., 2024b), MineDojo (Fan et al., 2022)
@@
TongUI (Zhang et al., 2025a), BEARCUBS (Song et al., 2025c)
@@
5.1 Multimodal Understanding
5.1.1 Visual-Centric Understanding
General Visual Understanding
Document, Chart, and OCR Visual Understanding
Document, chart, and OCR-based VQA datasets form a specialized domain focusing on understanding structured visual information that includes textual elements. Document VQA, exemplified by DocVQA (Mathew et al., 2021), targets document understanding, requiring models to locate and interpret text within documents to answer questions. Chart VQA, such as DVQA (Kafle et al., 2018), focuses on interpreting visual data representations, including bar charts, line graphs, and pie charts, testing the model’s ability to understand these structures. OCR-VQA datasets like TextVQA (Singh et al., 2019) and OCR-VQA (Mishra et al., 2019) emphasize reading and reasoning about text embedded within natural images. These datasets share several distinctive characteristics: 1) the critical integration of OCR with visual understanding, 2) multi-step reasoning that combines both textual and visual elements, and 3) domain-specific knowledge about document structures, chart conventions, or text layouts. Unlike general VQA datasets, these collections heavily emphasize the interplay between visual and textual content, requiring models to bridge modalities in more structured contexts. Additionally, datasets like AI2D (Hiippala et al., 2021) focus on diagrams and structured visual representations, enhancing reasoning over graphical content.
Multilingual Visual Understanding
Multilingual visual understanding datasets cater to the increasing demand for language diversity in multimodal systems. Datasets like CMMLU (Li et al., 2024f), C-Eval (Huang et al., 2023c), Exams-v (Das et al., 2024), M3exam (Zhang et al., 2023e), VideoVista-CulturalLingo (Chen et al., 2025e), and MTVQA (Tang et al., 2024) cover beyond English-centric VQA systems. These datasets are characterized by: 1) integration of questions and annotations in multiple languages, covering various language families, 2) testing visual understanding and linguistic capabilities across different cultural contexts, and 3) requiring models to understand visual concepts that may have specific cultural interpretations or references. Unlike single-language VQA datasets, these multilingual datasets evaluate and enhance the cross-lingual transfer abilities of MLLMs.
Video Understanding
@@ 5.1.1 Visual-Centric Understanding
Integrated evaluation benchmarks, such as MMBench (Liu et al., 2024f), Seed-Bench (Li et al., 2023c), and MME-RealWorld (Zhang et al., 2024f), have emerged to provide a more holistic evaluation of existing multimodal models. These benchmarks test how well models integrate visual and linguistic understanding in real-world scenarios, including 1) multidimensional evaluation frameworks that assess various aspects of visual understanding, from perception to reasoning and knowledge integration, 2) carefully designed questions aimed at exploring specific abilities and identifying weaknesses, and 3) standardized evaluation pipelines for fair comparison across models. Unlike early task-specific datasets, these benchmarks offer a comprehensive measure of models’ overall capabilities.
@@ 5.1.2 Audio-Centric Understanding
As Large Audio-Language Models (LALMs) continue to evolve, more models now possess the ability to understand both speech and diverse sounds. Consequently, researchers are proposing new evaluation benchmarks to comprehensively assess models’ audio understanding capabilities. VoiceBench (Chen et al., 2024h) focuses on models’ ability to understand speech in varied contexts, including evaluations of basic capabilities, colloquial expressions, and performance in noisy environments. AudioBench (Wang et al., 2024a) integrates diverse speech tasks (e.g., Automatic Speech Recognition, Speech Question Answering), sound tasks (e.g., Audio Captioning, Audio Question Answering), and tasks related to human voices (e.g., accent, age, and gender). Air-Bench (Yang et al., 2024d) and MMAU (Sakshi et al., 2024) expand upon this by including music tasks in their evaluations. SD-eval (Ao et al., 2024) combines speech tasks with environmental sound tasks, enabling models to understand complex, mixed audio scenarios. These benchmarks not only incorporate earlier evaluation methods but also provide a more comprehensive framework for assessing speech understanding across a wide range of real-world applications.
@@ 5.2.1 Cross-modal Generation
The field of text-to-image generation (T2I) has seen significant advancements, driven by diverse datasets and benchmarks tailored to tasks such as text-to-image generation, editing, and conditional generation.
For text-to-image generation, datasets like MSCOCO (30K) (Lin et al., 2014a), CC12M (Changpinyo et al., 2021), and Flickr30k (Plummer et al., 2017) offer large-scale, general-purpose image-text pairs, emphasizing everyday scenes and objects. In contrast, datasets like RedCaps (Desai et al., 2021) and COMMONPOOL (Gadre et al., 2023) introduce more complex text descriptions and higher-resolution images. Benchmarks such as GenEval (Ghosh et al., 2023) and ELLA (Hu et al., 2024a) focus on evaluating text-to-image alignment, assessing how accurately the generated images match the textual descriptions. Meanwhile, GenAI-Bench (Li et al., 2024a) and T2I-CompBench++ (Huang et al., 2023a) emphasize the handling of complex prompts and object interactions, highlighting the need for effective compositional generation and improved semantic alignment.
@@
For conditional text-to-image generation, datasets like ADE20K (Zhou et al., 2016) and CocoStuff (Caesar et al., 2016) offer detailed segmentation maps and scene parsing annotations, enabling models to generate images with specific scene structures. UniControl (Qin et al., 2023) introduces more comprehensive data, requiring models to handle multiple conditional inputs simultaneously. Benchmarks like UniCombine (Wang et al., 2025c) focus on evaluating instruction execution completeness, visual coherence, and consistency with constraints.
@@
In text-to-video generation, high-quality datasets and comprehensive benchmarks are critical for advancing research. Datasets like VidGen-1M (Tan et al., 2024b), OpenVid-1M (Nan et al., 2024), and VidProM (Wang & Yang, 2024) cover a wide range of video content and corresponding descriptive texts. Benchmarking tools such as AIGCBench (Fan et al., 2019), EvalCrafter (Liu et al., 2024e), and VBench (Huang et al., 2024a) evaluate models across various metrics like relevance, coherence, and visual quality. Specialized benchmarks like VideoScore (He et al., 2024), WorldSimBench (Qin et al., 2024), and WorldScore (Duan et al., 2025) expand evaluation to cover video quality and real-world accuracy, with VideoScore assessing user satisfaction.

### v2

@@ 4.3 Technical Prospects
Takeaways: Native Large Multimodal Reasoning Model (LMRMs) In this section, we examined the latest large multimodal model (e.g., O3 and O4-mini) and their performance on challenging tasks and benchmarks. We then presented the future trajectory for native multimodal large models in terms of capability scope and level, including omnimodal perception and understanding, multimodal interactive generative reasoning, and intelligent agent behaviour. To realize this vision, we discussed approaches related to unified perception, learning methods, and data synthesis. We hope that native LMRMs will achieve comprehensive perception, precise understanding, and deep reasoning as a paradigm shift in machine intelligence.
@@ 5 Dataset and Benchmark
Table 14: Overview of Multimodal Benchmarks and Training Datasets, categorized by task: Understanding (Visual-centric, Audio-centric) and Generation (Cross-modal, Joint Multimodal). These benchmarks often require short or long reasoning for successful task completion, e.g., challenging visual and audio generation.
@@
Seed-Bench (Li et al., 2023b), MME-RealWorld (Zhang et al., 2024f) MS-COCO (Lin et al., 2014), Virpt (Yang et al., 2024a)
@@
MMStar (Chen et al., 2024f), TRIG-Bench (Li et al., 2025e) VideoCC3M (Nagrani et al., 2022), FILIP (Yao et al., 2021)
@@
TikTalkCoref (Li et al., 2025i), AgMMU (Gauba et al., 2025) TikTalkCoref (Li et al., 2025i), MRES-32M (Liu et al., 2025c)
@@
RefCOCOm (Liu et al., 2025c), EmotionHallucer (Xing et al., 2025a)
SBVQA (Alasmary & Al-Ahmadi, 2023), H2VU-Benchmark (Wu et al., 2025e)
4D-Bench (Zhu et al., 2025e), V2P-Bench (Zhao et al., 2025f)
RSMMVP (Adejumo et al., 2025), HIS-Bench (Zhao et al., 2025c)
MMLA (Zhang et al., 2025d), SARLANG-1M (Wei et al., 2025e)
CasualVQA (Foss et al., 2025), SeriesBench (Zhang et al., 2025c)
WebUIBench (Lin et al., 2025c), MLLM-CL (Zhao et al., 2025b)
MERIT (Chow et al., 2025a), A4Bench (Wang et al., 2025o)
VLM@school (Peinl & Tischler, 2025), UnLOK-VQA (Patil et al., 2025)
DocMark (Xiao et al., 2025), KnowRecall and VisRecall (Wang et al., 2025g)
@@
Air-Bench (Yang et al., 2024e), MMAU (Sakshi et al., 2024) CoVoST2 (Wang et al., 2020), SIFT-50M (Pandey et al., 2025)
@@
MusicNet (Thickstun et al., 2017), AVE-PM (Liu et al., 2025f) ClothoAQA (Lipping et al., 2022), MusicNet (Thickstun et al., 2017)
ACVUBench (Yang et al., 2025h) NSynth (Engel et al., 2017), MusicCaps (Agostinelli et al., 2023)
Multimodal Generation Cross-modal Generation GenEval (Ghosh et al., 2023), T2I-CompBench++ (Huang et al., 2025a) MS-COCO (Lin et al., 2014), Flickr30k (Plummer et al., 2017)
@@
MJHQ-30K (Li et al., 2024d), VBench 2.0  (Zheng et al., 2025b) UniSim (Yang et al., 2024f), VidProM (Wang & Yang, 2024)
@@
LMM4LMM (Wang et al., 2025m), RISEBench (Zhao et al., 2025e) HQ-Edit (Hui et al., 2024), InstructPix2Pix (Brooks et al., 2023)
WikiVideo (Martin et al., 2025), OmniDiff (Liu et al., 2025g) MagicBrush (Zhang et al., 2023d)
@@
CompAlign (Wan & Chang, 2025), GODBench (Lei et al., 2025)
5.1 Multimodal Understanding
5.1.1 Visual-Centric Understanding
General Visual Understanding
To further stress compositional and spatial reasoning capabilities, MMSI-Bench introduces a spatial reasoning benchmark focused on evaluating multi-modal understanding of object configurations and spatial relationships (Yang et al., 2025e). WikiMixQA pushes models to integrate and synthesize multimodal information from textual and visual sources to answer complex queries that require cross-source reasoning (Foroutan et al., 2025). In the mathematical domain, VideoMathQA presents expert-curated, video-based math problems across three reasoning categories—direct problem solving, conceptual transfer, and structured explanation—mirroring real-world educational settings (Rasheed et al., 2025).
Document, Chart, and OCR Visual Understanding
Document, chart, and OCR-based VQA datasets form a specialized domain focusing on understanding structured visual information that includes textual elements. Document VQA, exemplified by DocVQA (Mathew et al., 2021), targets document understanding, requiring models to locate and interpret text within documents to answer questions. Chart VQA, such as DVQA (Kafle et al., 2018), focuses on interpreting visual data representations, including bar charts, line graphs, and pie charts, testing the model’s ability to understand these structures. OCR-VQA datasets like TextVQA (Singh et al., 2019) and OCR-VQA (Mishra et al., 2019) emphasize reading and reasoning about text embedded within natural images. These datasets share several distinctive characteristics: 1) the critical integration of OCR with visual understanding, 2) multi-step reasoning that combines both textual and visual elements, and 3) domain-specific knowledge about document structures, chart conventions, or text layouts. Unlike general VQA datasets, these collections heavily emphasize the interplay between visual and textual content, requiring models to bridge modalities in more structured contexts. Additionally, datasets like AI2D (Hiippala et al., 2021) focus on diagrams and structured visual representations, enhancing reasoning over graphical content. WebUIBench introduces a structured benchmark for planning and reasoning within digital interface environments. It evaluates MLLMs across four dimensions—WebUI perception, HTML programming, WebUI-HTML understanding, and UI-to-code generation—making it a comprehensive testbed for web interface grounding and control (Lin et al., 2025c).
Multilingual Visual Understanding
Multilingual visual understanding datasets cater to the increasing demand for language diversity in multimodal systems. Datasets like CMMLU (Li et al., 2024f), C-Eval (Huang et al., 2023c), Exams-v (Das et al., 2024), M3exam (Zhang et al., 2023e), VideoVista-CulturalLingo (Chen et al., 2025h), and MTVQA (Tang et al., 2024) cover beyond English-centric VQA systems. These datasets are characterized by: 1) integration of questions and annotations in multiple languages, covering various language families, 2) testing visual understanding and linguistic capabilities across different cultural contexts, and 3) requiring models to understand visual concepts that may have specific cultural interpretations or references. Unlike single-language VQA datasets, these multilingual datasets evaluate and enhance the cross-lingual transfer abilities of MLLMs.
New benchmarks also expand the scope of visual understanding to multilingual and real-world contexts. CasualVQA emphasizes causal and contextual reasoning via questions grounded in everyday scenarios, aiming to test models’ deeper comprehension of real-world phenomena (Foss et al., 2025). Meanwhile, VLM@school evaluates models’ ability to integrate visual reasoning with curriculum-based subject knowledge in German, supporting multilingual educational applications (Peinl & Tischler, 2025).
Video Understanding
Table 15: Overview of Multimodal Benchmarks and Training Datasets, categorized by task: Reasoning (General Visual, Domain-Specific) and Planning (GUI, Embodied & Simulated Environments). These benchmarks often require short or long reasoning for successful task completion, e.g., challenging visual and audio generation.
Ability Task Benchmark Dataset
@@ 5.1.1 Visual-Centric Understanding
PhysBench (Chow et al., 2025b), MMBench (Liu et al., 2024f) MMPR (Wang et al., 2024g), ChartQA (Masry et al., 2022)
@@
VCRBench (Qi et al., 2025), VisualPuzzles (Song et al., 2025f) ShareGPT4oReasoning (Zhang et al., 2024e), R1-Onevision (Yang et al., 2025g)
IV-Bench (Ma et al., 2025a), VisuLogic (Xu et al., 2025f) Video-R1-data (Feng et al., 2025b), Visual-CoT (Shao et al., 2024)
FG-BMK (Yu et al., 2025a), Video-MMLU (Song et al., 2025a)
@@
HAVEN (Gao et al., 2025a), MAGIC-VQA (Yang et al., 2025d)
@@
FortisAVQA (Ma et al., 2025b), VideoVista-CulturalLingo (Chen et al., 2025h)
CoMT (Cheng et al., 2025b), CCHall (Zhang et al., 2025o)
Wethink-dataset (Yang et al., 2025b), MMMG (Yao et al., 2025c)
Do You See Me (Kanade & Ganu, 2025), SpaCE-10 (Gong et al., 2025)
GaslightingBench-R (Zhu et al., 2025a), MultiNet (Guruprasad et al., 2025)
MANBench (Zhou et al., 2025a), VGR -SFT (Wang et al., 2025k)
TempVS benchmark (Song et al., 2025d), SFE benchmark (Zhou et al., 2025d)
HSSBench (Kang et al., 2025), Argus Inspection (Yao et al., 2025d)
Chemtable (Zhou et al., 2025c), KokushiMD-10 (Liu et al., 2025d)
@@
ScienceQA (Lu et al., 2022), ChartQA (Masry et al., 2022) SpaceR-151k (Ouyang et al., 2025)
@@
VCBench (Wang et al., 2025w), VCT (Götting et al., 2025)
3MDBench (Sviridov et al., 2025), PuzzleBench (Zhang et al., 2025q)
ColorBench (Liang et al., 2025b), VisualPuzzles (Song et al., 2025f)
@@
EIBench (Lin et al., 2025b), XLRS-Bench (Wang et al., 2025e)
STI-Bench (Li et al., 2025l), EgoToM (Li et al., 2025p)
DomainCQA (Zhong et al., 2025b), MM-IQ (Cai et al., 2025)
MMCR-Bench (Yan et al., 2025a), Misleading ChartQA (Chen et al., 2025m)
@@
MicroVQA (Burgess et al., 2025), MPBench (Xu et al., 2025h)
Open3DVQA (Zhan et al., 2025), ProBench (Yang et al., 2025f)
Chart-HQA (Chen et al., 2025f), MMSciBench (Ye et al., 2025b)
CharXiv(Wang et al., 2024n), SPAR-Bench(Zhang et al., 2025f)
VSI-Bench(Yang et al., 2024b), STI-Bench(Li et al., 2025l)
VRBench (Yu et al., 2025b), MATP-BENCH (He et al., 2025)
MMRefine (Paik et al., 2025), MMR-V (Zhu et al., 2025b)
WikiMixQA (Foroutan et al., 2025), SciVer (Wang et al., 2025c)
MultiFinBen (Peng et al., 2025), RealHiTBench (Wu et al., 2025d)
VideoMathQA (Rasheed et al., 2025), GDI-Bench (Li et al., 2025g)
PhysicsArena (Dai et al., 2025b), Spoken-MQA (Wei et al., 2025a)
ReasonSeg-Diff (Wang et al., 2025q), VF-Eval (Song et al., 2025c)
ChartMind (Wei et al., 2025c)
@@
SWE-bench M (Yang et al., 2024c), MineDojo (Fan et al., 2022)
@@
TongUI (Zhang et al., 2025a), BEARCUBS (Song et al., 2025e)
MCA-Bench (Wu et al., 2025h), ThinkGeo (Shabbir et al., 2025)
@@
AttachSecure-Bench (Wang et al., 2025a)
@@
Integrated evaluation benchmarks, such as MMBench (Liu et al., 2024f), Seed-Bench (Li et al., 2023b), and MME-RealWorld (Zhang et al., 2024f), OmniBench  (Li et al., 2024j). This benchmark enables unified evaluation across reasoning, planning, and generation tasks, and supports platform-agnostic, scalable testing of MLLM capabilities. have emerged to provide a more holistic evaluation of existing multimodal models. These benchmarks test how well models integrate visual and linguistic understanding in real-world scenarios, including 1) multidimensional evaluation frameworks that assess various aspects of visual understanding, from perception to reasoning and knowledge integration, 2) carefully designed questions aimed at exploring specific abilities and identifying weaknesses, and 3) standardized evaluation pipelines for fair comparison across models. Unlike early task-specific datasets, these benchmarks offer a comprehensive measure of models’ overall capabilities.
@@ 5.1.2 Audio-Centric Understanding
As Large Audio-Language Models (LALMs) continue to evolve, more models now possess the ability to understand both speech and diverse sounds. Consequently, researchers are proposing new evaluation benchmarks to comprehensively assess models’ audio understanding capabilities. VoiceBench (Chen et al., 2024h) focuses on models’ ability to understand speech in varied contexts, including evaluations of basic capabilities, colloquial expressions, and performance in noisy environments. AudioBench (Wang et al., 2024a) integrates diverse speech tasks (e.g., Automatic Speech Recognition, Speech Question Answering), sound tasks (e.g., Audio Captioning, Audio Question Answering), and tasks related to human voices (e.g., accent, age, and gender). Air-Bench (Yang et al., 2024e) and MMAU (Sakshi et al., 2024) expand upon this by including music tasks in their evaluations. SD-eval (Ao et al., 2024) combines speech tasks with environmental sound tasks, enabling models to understand complex, mixed audio scenarios. These benchmarks not only incorporate earlier evaluation methods but also provide a more comprehensive framework for assessing speech understanding across a wide range of real-world applications.
@@ 5.2.1 Cross-modal Generation
The field of text-to-image generation (T2I) has seen significant advancements, driven by diverse datasets and benchmarks tailored to tasks such as text-to-image generation, editing, and conditional generation. For text-to-image generation, datasets like MSCOCO (30K) (Lin et al., 2014), CC12M (Changpinyo et al., 2021), and Flickr30k (Plummer et al., 2017) offer large-scale, general-purpose image-text pairs, emphasizing everyday scenes and objects. In contrast, datasets like RedCaps (Desai et al., 2021) and COMMONPOOL (Gadre et al., 2023) introduce more complex text descriptions and higher-resolution images. Benchmarks such as GenEval (Ghosh et al., 2023) and ELLA (Hu et al., 2024a) focus on evaluating text-to-image alignment, assessing how accurately the generated images match the textual descriptions. Meanwhile, GenAI-Bench (Li et al., 2024a) and T2I-CompBench++ (Huang et al., 2023a) emphasize the handling of complex prompts and object interactions, highlighting the need for effective compositional generation and improved semantic alignment.
@@
For conditional text-to-image generation, datasets like ADE20K (Zhou et al., 2016) and CocoStuff (Caesar et al., 2016) offer detailed segmentation maps and scene parsing annotations, enabling models to generate images with specific scene structures. UniControl (Qin et al., 2023) introduces more comprehensive data, requiring models to handle multiple conditional inputs simultaneously. Benchmarks like UniCombine (Wang et al., 2025h) focus on evaluating instruction execution completeness, visual coherence, and consistency with constraints.
@@
In text-to-video generation, high-quality datasets and comprehensive benchmarks are critical for advancing research. Datasets like VidGen-1M (Tan et al., 2024b), OpenVid-1M (Nan et al., 2024), and VidProM (Wang & Yang, 2024) cover a wide range of video content and corresponding descriptive texts. Benchmarking tools such as AIGCBench (Fan et al., 2019), EvalCrafter (Liu et al., 2024e), and VBench (Huang et al., 2024a) evaluate models across various metrics like relevance, coherence, and visual quality. Specialized benchmarks like VideoScore (He et al., 2024), WorldSimBench (Qin et al., 2024), and WorldScore (Duan et al., 2025b) expand evaluation to cover video quality and real-world accuracy, with VideoScore assessing user satisfaction.